The three times of the, sum of ten and half of a number is sum of 6 and, a number decreased by 5
Answer by swincher4391(1107)

Not sure what the first part is all about, but I will read it as "sum of ten and half of a number is sum of 6 and, a number decreased by 5"
10 + n/2 = 6 + n-5
10+ n/2 = n + 1
20 + n = 2n + 2
-n = -18
n = 18
Check:
10 + 9 = 6 + 13